Write the possible term symbols for

(a) Li: [He] 2s1

(b) Na: [Ne] 3p1
Answer #1

(a) Li [He]2S1 2SH = 2*1/2 + 1 = 2 L = 0 = > S J = LiS = (1/2) = 1/2 Term symbol = 2S1/2

(b)The electronic configuration of Na is [Ne] 3s¹

So,for this given part the term symbol is not possible because the the configuration is wrong.

Hope you like the solution.If you have any further queries,you can ask in the comment box and i will definitely have them solved for you.Please hit the like button,please.